### v1

# UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality

Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human values. However, existing approaches struggle to capture the multi-dimensional, distributional nuances of human preferences. Methods such as RiC that directly inject raw reward values into prompts face significant numerical sensitivity issues—for instance, LLMs may fail to distinguish between 9.11 and 9.8—while alternatives like MORLHF, Rewarded Soups, and MODPO incur high computational costs by training multiple models. In this work, we introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework that overcomes these limitations. Our approach leverages a diverse set of strictly increasing, non-linear utility functions to transform user-specified preferences into symbolic tokens, which are then used to condition a single LLM. This design not only mitigates numerical reasoning challenges but also substantially reduces training overhead, yielding models that achieve superior Pareto fronts and robust alignment across complex reward dimensions.
UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality
Zelei Chengthanks: Equal contribution. Northwestern University Evanston, USA                        Xin-Qiang Cai11footnotemark: 1 RIKEN-AIP Tokyo, Japan                        Yuting Tang The University of Tokyo Tokyo, Japan
Pushi Zhang Microsoft Research Asia Beijing, China                        Boming Yang The University of Tokyo Tokyo, Japan                        Xinyu Xing Northwestern University Evanston, USA
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a powerful paradigm for aligning artificial agents with human values by incorporating user-provided feedback into the training process Ouyang et al. (2022); Bai et al. (2022); Stiennon et al. (2020). Despite its success, conventional RLHF approaches typically rely on scalar reward functions and struggle to capture the full spectrum of human preferences, which are inherently multi-dimensional and distributional in nature. For example, methods like RiC (Yang et al., 2024c) directly input raw reward values into the prompts. However, LLMs are known to have limited numerical reasoning capabilities, often failing to distinguish between numerical inputs (e.g., which is bigger? 9.11 versus 9.8 (Yang et al., 2024b)). This numerical insensitivity can lead to suboptimal alignment when dealing with nuanced feedback that spans multiple reward dimensions.
Concurrently, alternative strategies such as MORLHF (Li et al., 2020), Rewarded Soups (Rame et al., 2023), and MODPO (Zhou et al., 2024) attempt to address the complexity of multi-objective alignment by training multiple models to handle different reward combinations Roijers et al. (2013). While these methods provide a richer representation of human preferences, they are computationally expensive and scale poorly with the number of reward dimensions, rendering them impractical for large-scale applications. The high training cost not only limits their real-world deployment but also hinders rapid experimentation and adaptation in dynamic environments.
Figure 1: Illustration of the training processes for existing multi-objective alignment methods compared to our proposed UC-MOA. While methods such as RiC and Rewarded Soups directly inject raw reward values (which are prone to numerical errors) or require multiple models to capture diverse preferences, UC-MOA employs a single model conditioned on a symbolic token derived from a diverse set of strictly increasing utility functions.
To overcome these challenges, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA) for Distributional Pareto-Optimality. Our framework builds upon recent advances in multi-objective reinforcement learning, where the concept of Pareto-optimality is extended to distributional Pareto-optimality—a notion that prioritizes the entire return distribution over mere expectations Levy (1992); Shaked and Shanthikumar (2007). The difference of our method compared with previous methods are illustrated in Figure 1. Central to our approach is the use of a diverse set of strictly increasing, non-linear utility functions, parameterized via monotonic neural networks Amos et al. (2017); Dugas et al. (2009). These utility functions are trained on offline data with human feedback, enabling them to capture not only the expected values but also higher-order statistical properties such as variance and skewness. By converting user preferences into target rewards through a linear mapping and subsequently conditioning the LLM on a symbolic token (i.e., <max_utility_index> concatenated with the optimal utility index), UC-MOA effectively circumvents the numerical sensitivity issues that plague raw reward injection.
In our framework, the use of utility functions not only resolves the numerical reasoning challenges of previous methods but also introduces a principled mechanism to attain distributional Pareto-optimality. By ranking the outputs of these utility functions based on their percentile scores, we dynamically select the most appropriate utility perspective during training—ensuring that the resulting policy optimally balances conflicting objectives in a distributional sense. This strategy enables a single model to be conditioned on complex, multi-dimensional reward profiles, thereby dramatically reducing the training cost compared to multi-model approaches while achieving superior Pareto fronts. Our empirical evaluations demonstrate that UC-MOA robustly aligns LLM outputs with nuanced human preferences across various tasks, offering a scalable and effective solution for real-world multi-objective alignment.
2 Related Work
2.1 Large Language Model Alignment
Aligning LLMs with human preferences is a critical challenge, and RLHF has emerged as a cornerstone technique in this domain. Initially popularized by Christiano et al. (2017) using a Proximal Policy Optimization (PPO)-based framework, RLHF has played a key role in developing advanced LLMs such as ChatGPT and Claude (Bai et al., 2022; Ouyang et al., 2022). However, scaling PPO for large models introduces significant efficiency challenges (Choshen et al., 2020; Engstrom et al., 2020), motivating the development of more scalable alternatives. Recent methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and the Reward-Ranked Fine-Tuning (RAFT) algorithm (Dong et al., 2023) specifically address alignment challenges in LLMs. DPO directly optimizes preference datasets without relying on explicit reward models, and its online iterative variants further improve exploration and robustness (Xiong et al., 2024; Liu et al., 2024). In parallel, RAFT has been successfully applied in LLMs such as LLaMA3 (Dubey et al., 2024) and Qwen2 (Yang et al., 2024a) by sampling multiple responses per prompt, ranking them via a reward model, and fine-tuning the model on the top-ranked outputs—a strategy shown to enhance reasoning capabilities (Aksitov et al., 2023; Havrilla et al., 2024).
2.2 Multi-objective Alignment
In contrast to traditional RLHF, which optimizes a single reward function, multi-objective alignment aims to balance multiple criteria simultaneously, such as harmlessness, humor, and helpfulness. This approach is crucial for real-world applications where conflicting user preferences and diverse objectives must be integrated into a model’s behavior.
Standard MORLHF (Wang et al., 2024) aggregates multiple reward signals via a linear scalarization strategy, maximizing a weighted sum of scores. However, because this method fixes the weight vector during training, it inherently supports only a single, static user preference. To overcome this limitation and better reflect the nuances of multi-objective alignment, Rewarded Soups (Rame et al., 2023) trains $N$ separate policies—each dedicated to a specific objective—and then linearly interpolates these policies. This process produces a continuous spectrum of models that can be tuned to diverse reward configurations, though it introduces scalability challenges due to the cost of training multiple policies.
Alternatively, RiC Yang et al. (2024c) adapts to different user preferences by conditioning model outputs on multiple reward signals embedded within the prompt. While this method offers flexibility, it still encounters the inherent numerical reasoning limitations of large language models, which struggle to parse and integrate raw numerical reward signals during generation.
Our proposed technique tackles the core challenges of multi-objective alignment by pre-training a diverse set of strictly increasing, non-linear utility functions. We then inject a symbolic token into the prompt that encodes the maximum utility index, enabling the model to indirectly condition its output on the reward structure without relying on precise numerical calculations. This design directly addresses the complexity of multi-objective alignment by allowing the model to seamlessly balance and integrate multiple objectives according to varied user preferences.
2.3 Distributional Reinforcement Learning.
Traditional RL focuses on estimating the expected return of a policy, but distributional reinforcement learning (DRL) extends this framework by modeling the entire distribution of returns rather than just their expectation Bellemare et al. (2017); Zhang et al. (2021). This richer representation of uncertainty has been shown to enhance both learning efficiency and policy quality in various single-objective RL tasks Dabney et al. (2018); Barth-Maron et al. (2018).
Several key algorithms have been developed in this area. Categorical DQN Bellemare et al. (2017) approximates the return distribution using a fixed set of discrete atoms, while Quantile Regression DQN Dabney et al. (2018) models it through quantile regression. Another approach, Distributional MPO Barth-Maron et al. (2018), enhances Maximum a Posteriori Policy Optimization by incorporating distributional learning, leading to improved exploration and stability.
Despite these successes, extending distributional RL to MORL remains an open challenge. The presence of multiple, often conflicting, objectives increases the complexity of reasoning about the full return distribution, as well as how preferences should be expressed over these distributions. Recent work by Cai et al. (2023) introduces Distributional Pareto-Optimal Multi-Objective Reinforcement Learning (DPMORL), which seeks to learn a set of Distributional Pareto-Optimal policies. They show that for any strictly increasing utility function, the policy maximizing expected utility corresponds to a Distributional Pareto-Optimal policy. This result suggests a principled way to incorporate distributional reasoning into MORL while preserving Pareto efficiency. Building on this theoretical foundation, our method distinguishes itself by generating a rich ensemble of non-linear utility functions using monotonic neural networks. By incorporating a symbolic selection mechanism—wherein the maximum utility index is determined and appended to the prompt—we effectively decouple the numerical complexities from the LLM’s generation process. This allows our framework to seamlessly integrate distributional considerations into the multi-objective setting while preserving the desirable properties of Pareto efficiency.
3 Proposed Technique
In this work, we propose a novel framework that integrates distributional utility modeling with RLHF. Recent work such as RiC (Yang et al., 2024d) and Rewarded Soups (Rame et al., 2023) directly appends raw reward values to the prompts. However, as argued in Section 1, LLMs exhibit limited capability in directly processing precise numerical inputs. Instead, we first train a diverse set of non-linear utility functions from an offline dataset and then introduce a symbolic token into the prompt. This token consists of the index corresponding to the maximum utility value, denoted as <max_utility_index>, concatenated with its associated utility index. During online fine-tuning, after generating a response, we determine the maximal utility index by ranking the outputs of each utility function based on their percentile scores, and then update the LLM via an online process that follows the RiC paradigm except for this novel selection mechanism. Figure 1 shows the framework of our method.
3.1 Utility Function Generation with Human Feedback
The first component of our algorithm is dedicated to constructing a rich set of candidate utility functions that will later guide the derivation of diverse optimal policies. Such diversity is crucial for addressing various distributional preferences and for adapting to different application scenarios. To achieve this, we (i) parameterize each candidate utility function using a monotonic neural network, and (ii) optimize a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities of these functions over a normalized domain.
Monotonic Neural Network Parameterization. We denote each candidate utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized input space. The network $g_{\psi_{i}}$ is constructed to be monotonic (i.e., non-decreasing) by enforcing non-negativity on its weight matrices and by employing non-decreasing activation functions. This design paradigm follows the approaches proposed in convex neural networks Amos et al. (2017) and monotonic architectures such as DPMORL Cai et al. (2023), which have been shown to approximate any multivariate non-decreasing function with arbitrarily small error Dugas et al. (2009). Implementation details are provided in Appendix B.
Diversity-Promoting Objective. To ensure that the candidate functions span a wide range of utility profiles, we introduce a loss function that encourages dissimilarity both in function values and in their gradients. For a given candidate $g_{\psi_{i}}$, we first define the value discrepancy loss as
$L_{\mathrm{val}}(\psi_{i})=\min_{j\neq i}\mathbb{E}_{\mathbf{z}\sim\mathcal{U}(\mathcal{Z})}\Bigl{[}g_{\psi_{i}}(\mathbf{z})-g_{\psi_{j}}(\mathbf{z})\Bigr{]}^{2},$ (1)
where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
$\begin{split}L_{\mathrm{grad}}&(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\\ &\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
The overall diversity loss for the candidate $g_{\psi_{i}}$ is then defined as a convex combination of these two components:
$L(\psi_{i})=\mu\,L_{\mathrm{val}}(\psi_{i})+(1-\mu)\,L_{\mathrm{grad}}(\psi_{i}),$ (3)
where the hyperparameter $\mu\in[0,1]$ balances the relative emphasis on value versus gradient discrepancies. Minimizing $L(\psi_{i})$ via gradient descent for each $i=1,\dots,M$ encourages the ensemble $\{g_{\psi_{i}}\}_{i=1}^{M}$ to cover a broad spectrum of utility profiles over $\mathcal{Z}$.
To satisfy the strict monotonicity condition required for ensuring Distributional Pareto-Optimality, we further modify each candidate by incorporating a small linear term. Specifically, we define the final utility function as
$\hat{g}_{\psi_{i}}(\mathbf{z})=g_{\psi_{i}}(\mathbf{z})+\frac{\epsilon}{K}\sum_{k=1}^{K}z_{k},$ (4)
where $\epsilon$ is a small positive constant (e.g., $\epsilon=0.01$ in our experiments). The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ are then employed as the strictly increasing utility functions for guiding policy optimization.
3.2 Utility Index Selection during Training
To robustly determine the optimal utility index during training, we employ a percentile ranking strategy in both the offline and online phases. For every sample with reward vector $\mathbf{z}$, each candidate utility function $\hat{g}_{\psi_{i}}$ produces a utility score. We then compute the percentile rank for the $i$-th utility function as
$\text{percentile}_{i}=\frac{\text{rank of }\hat{g}_{\psi_{i}}(\mathbf{z})}{N},$ (5)
where $N$ is the total number of samples, and the rank is determined by the position of $\hat{g}_{\psi_{i}}(\mathbf{z})$ in the sorted array of scores for that function. The percentile ranking provides a normalized measure of performance for each utility function, mitigating the effects of scale differences and ensuring that selection is robust across diverse reward distributions. During offline and online training, the utility index is then selected as the one with the highest percentile rank. We present a concrete example in Appendix B.2.
3.3 Offline Training
In the offline phase, our approach establishes a robust grounding of the language model in distributional reward signals by embedding an explicit symbolic indicator of the preferred utility function directly into the training prompt. For each training sample, denoted by $(x,y)$ with $x$ representing the input prompt and $y$ the corresponding response, we first compute a multi-objective return vector $\mathbf{z}(x,y)$ and then evaluate a suite of strictly increasing candidate utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ on $\mathbf{z}(x,y)$. The scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(x,y))$ for $i=1,\dots,M$ are then obtained, and for each function, we calculate its percentile rank over the training distribution. The maximum utility index is determined as the index with the highest percentile:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}\left\{\text{percentile}_{i}\right\}.$ (6)
We augment the original prompt $x$ by appending a specialized token that encodes this index; that is, the modified prompt is given by
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The model is then fine-tuned via a standard cross-entropy loss:
$\mathcal{L}_{\text{offline}}(\theta)=-\sum_{t}\log\pi_{\theta}(y_{t}\mid x^{\prime},y_{<t}),$
ensuring that the model learns to internalize the mapping between input contexts and their corresponding optimal distributional reward profiles. This training stage leverages supervised techniques exclusively, thereby avoiding the instability often associated with RL-based methods.
3.4 Online Fine-Tuning
To enhance the generalizability of the fine-tuned model, we incorporate additional online stages. In the online stage, the large language model generates additional candidate responses and we selectively choose those responses that exhibit superior distributional reward characteristics to fine-tune the model. Given an input prompt $x$, the model produces a response $y$ whose multi-objective return $\mathbf{z}(y)$ is computed and then evaluated by each candidate utility function, yielding scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z}(y))$ for $i=1,\dots,M$. For each candidate response, we compute the percentile rank of each utility score as described above, and the optimal utility index is selected as the one with the highest percentile as in Equation (5), resulting in an updated input
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*}.$
The response $y$ and its corresponding modified prompt $x^{\prime}$ are stored in an online training buffer for further model updates. Our online training process also incorporates a rejection sampling mechanism to filter out responses that do not exceed predetermined utility thresholds across multiple reward dimensions. The selected samples are then used to update the language model via the same reward-conditional supervised objective as in the offline training phase.
3.5 Inference Stage
At inference time, the user specifies a preference vector $\mathbf{w}=[w_{1},\ldots,w_{N}]$ that encodes the relative importance of each reward dimension $r_{1},\ldots,r_{N}$. To translate these preferences into a concrete reward profile, we apply a linear preference-to-reward mapping for each dimension:
$f_{i}(w_{i})=w_{i}\times(r_{i}^{\mathrm{max}}-r_{i}^{\mathrm{min}})+r_{i}^{\mathrm{min}},$
where $r_{i}^{\mathrm{max}}$ and $r_{i}^{\mathrm{min}}$ denote the maximum and minimum values of the $i$-th reward observed in the training dataset. This mapping yields a target reward vector $\mathbf{z}=[f_{1}(w_{1}),\dots,f_{N}(w_{N})]$ that accurately reflects the user’s preferences while normalizing the reward scales across dimensions.
Subsequently, the target reward vector $\mathbf{z}$ is input into each of the pre-trained, strictly increasing utility functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ to obtain scalar utility values $U_{i}=\hat{g}_{\psi_{i}}(\mathbf{z})$ for $i=1,\dots,M$. at inference time the reward distribution is more stable and the set of candidate responses is smaller; therefore, we directly use the argmax operation to select the index corresponding to the maximum expected utility:
$i^{*}=\arg\max_{i\in\{1,\dots,M\}}U_{i}.$ (7)
This index $i^{*}$ is encoded into the input prompt via a symbolic token—specifically, the prompt is augmented as
$x^{\prime}=\text{\#\#\# Prompt: }\{x\}\;\texttt{<max\_utility\_index>}+i^{*},$
thereby conditioning the language model to generate responses that are robustly aligned with the user’s specified multi-objective reward preferences while obviating the need for direct numerical processing.
4 Evaluation
In this section, we start with the experimental setup, including the tasks, baselines, and implementation. We then present our experiment design, followed by experiment results and analysis. We provide our code in the anonymous link https://anonymous.4open.science/r/UC-MOA-8D36.
4.1 Experiment Setup
Tasks. Following RiC (Yang et al., 2024c), we select two text generation tasks for our experiments, i.e., Helpful Assistant task (Bai et al., 2022) and the Reddit Summary task (Stiennon et al., 2020). For the Helpful Assistant task, we utilize three open-sourced reward models on Huggingface to measure three different objectives “harmless”, “helpful”, and “humor”. We focus on three multi-objective subtasks: (1) “harmless” vs “helpful”; (2) “humor” vs “helpful”; (3) “harmless” vs “humor”. Regarding the Reddit Summary task, we consider three reward models namely “summary”, “hallucinated”, and “faithful” to evaluate different perspectives of the summary quality. We focus on three multi-objective subtasks: (1) “summary” vs “faithful”; (2) “hallucinated” vs “faithful”; (3) “summary” vs “hallucinated”. We provide more details about the datasets and the reward models in Appendix A.
Baseline methods.We compare our method with two state-of-the-art multi-objective RLHF methods, Rewarded Soups, and Reward-in-Context (RiC). We also report the performance of the base model and the supervised fine-tuned (SFT) model.
Implementation. We implement the proposed method using PyTorch. We train the machine learning models on a server with 8 NVIDIA A100 80GB GPUs and 4TB memory for all methods. We select the LLaMA 2 7B-hf model as the base model for alignment following the setting of Yang et al. (2024c). We list the implementation details (e.g., network structure for the parameterized utility function, hyper-parameter choices) in Appendix B.
4.2 Experiment Design
We use the following experiments to evaluate the effectiveness and efficiency of our method, and other factors that influence the performance (e.g., alternative design choices, hyper-parameters).
Experiment I. To evaluate the effectiveness of our method, we compare the curves of the multi-dimensional average test rewards that correspond to the empirical Pareto fronts generated by different methods. A method demonstrating an outer curve indicates better performance, as it captures a wider range of trade-offs across objectives. We plot the Pareto fronts of all baseline methods across the Helpful Assistant task and the Reddit Summary task.
Experiment II. We analyze the computational efficiency of each method by recording the training time of each baseline method. We run experiments with NVIDIA A100 80GB GPUs and report the GPU hours utilized by each method.
Experiment III. To validate the necessity of the online stage, we perform an ablation study to compare the performance of pure offline training (“w/o online”), a single online iteration (“w/ online iter 1”), and two online iterations (“w/ online iter 2”).
Experiment IV. We assess the consistency of our method during the inference phase by measuring the alignment between the user-specified target objective and the reward outcomes obtained from the LLM’s generated responses. Each prompt provided to the LLM embeds a target objective requirement (e.g., in the form <max_utility_index> a, where the letter encodes the desired objective). During evaluation, the LLM is tasked with generating responses that adhere to these preset requirements. For each generated response, we compute a set of utility values using a suite of ten reward models. To account for scale differences across objectives, the computed reward values are first normalized on a per-utility basis. We then quantify consistency by comparing the expected utility and the value distribution of each utility function with the target objective extracted from the prompt.
Experiment V. To investigate how the number of utility functions impacts the performance of our method, we vary the number of utility functions from {5,10,15} and test the performance of our method with different numbers of utility functions.
Experiment VI. We investigate an alternative design choice of the utility functions, i.e., linear functions. The linear function is designed in the form of $w^{T}x$ where $||w||_{2}=1$. To construct $N$ linear utility functions, we uniformly sample $N$ points from the unit circle and set $w$ accordingly. We plot the Pareto fronts using linear utility functions and our method across all tasks.
Experiment VII. We also investigate two metrics developed specifically for assessing distributional preference, i.e., Constraint Satisfaction and Variance Objective as in previous distributional studies (Cai et al., 2023). They are designed to underscore the optimality of multivariate distributions associated with LLMs. In terms of Constraint Satisfaction, we randomly generate $M=100$ constraints. The Constraint Satisfaction metric is then computed by considering the highest probability within the policy set that satisfies each individual constraint. The Variance Objective metric, on the other hand, involves generating $M=100$ random linear weights. These weights are applied to both the expected returns and the standard deviations of returns in each dimension. This objective encourages attaining greater expected returns while simultaneously increasing variance, thereby catering to dynamic preferences. Further details about the implementation of these evaluation metrics are provided in Appendix B.4.
4.3 Experiment Results.
Performance comparison. We compare the multi-dimensional average test rewards obtained from the empirical Pareto fronts generated by different methods on two distinct tasks: the Helpful Assistant task and the Reddit Summary task. We show the results of “harmless” vs. “humor” task and “deberta” vs. “faithful” in Figure 2 while leaving the rest of the results in Appendix C.1. We observe that our method consistently produces an outer Pareto curve that dominates those of Rewarded Soups and RiC, indicating its superior ability to capture a broader spectrum of trade-offs.
Figure 2: Pareto fronts of “harmless” vs “humor” and “deberta” vs. “faithful” tasks. Each marker on the plotted curves represents the average test rewards corresponding to a particular user preference.
Computational cost. We measure the time complexity of all methods in the two-objective Helpful Assistant task and the two-objective Reddit Summary task using NVIDIA A100 80GB GPUs. We report the time cost (in the form of GPU hours) in Table 1 and Table 2. We observe that Rewarded Soups consistently consumes the most GPU hours across both tasks, while our method (w/ online iter2) achieves notably lower costs. Meanwhile, RiC requires more computing time than our method, though it remains more cost-friendly than Rewarded Soups. These findings demonstrate that our approach not only delivers strong performance (as shown in previous experiments) but also offers a clear advantage in terms of computational efficiency.
  Method GPU hours
  Rewarded Soups 69.5
RiC 20.3
Ours w/o online 8.9
Ours w/ online iter1 10.1
Ours w/ online iter2 11.3
Table 1: Comparison of GPU hours in the Helpful Assistant task, where the number of preference $M=10$, number of rewards $N=2$.
  Method GPU hours
  Rewarded Soups 15.9
RiC 5.7
Ours w/o online 2.0
Ours w/ online iter1 2.4
Ours w/ online iter2 2.8
Table 2: Comparison of GPU hours in the Reddit Summary task, where the number of preference $M=10$, number of rewards $N=2$.
Necessity of online stage. We evaluate the impact of the online stage by comparing models trained purely offline (“w/o online”) with those undergoing one (“w/ online iter 1”) and two online iterations (“w/ online iter 2”) across all tasks. We present the performance results for “harmless” vs. “humor” and “summary” vs. “faithful” tasks in Figure 3 while leaving the rest of the results in Appendix C.2. We observe that models incorporating online iterations consistently outperform the offline-only baseline. Notably, even a single online iteration yields significant improvements, with further performance gains observed when two iterations are applied, thereby underscoring the necessity of the online stage in our training framework. The online stage is important because it enables the model to generalize over unseen samples in the offline dataset, and ultimately enhancing its ability to generate appropriate responses corresponding to the required utility function index.
Figure 3: Ablation study of the online stage on “harmless” vs. “humor” task and “summary” vs. “faithful” task.
Consistency. We conduct a case study on the “harmless” vs. “humor” task, appending the requirement “<max_utility_idx> i” to each prompt to encourage the model to follow the first preference (i.e., maximizing 9th utility function). We then use reward models to evaluate the generated responses. We provide a detailed analysis in Appendix C.3 to present the distribution of normalized utility values for each utility function over our evaluation data. Notably, the 9th utility function achieves the highest mean normalized value, indicating that the fine-tuned model aligns its responses with the specified preference requirement. This demonstrates the model’s ability to selectively prioritize and adhere to user-designated objectives.
Number of utility functions. We vary the number of utility functions over the set 5, 10, 15 and compare the resulting empirical Pareto fronts in Figure 4. We leave the results of the remaining four tasks in Appendix C.4. With 5 utility functions, the generated Pareto front is slightly inferior to those produced with 10 and 15 utility functions. Specifically, the front with 5 utility functions captures a narrower range of optimal trade-offs across the objectives. However, when the number of utility functions is increased to 10 or 15, the resulting Pareto fronts exhibit comparable performance. Increasing the number of utility functions beyond 10 does not yield substantial gains. These results demonstrate that our method is insensitive to the precise number of utility functions, provided that a sufficient number (around 10) is used to effectively cover the preference space. In practical terms, employing 10 utility functions strikes a good balance between performance and computational efficiency, as further increases to 15 do not produce notable improvements.
Figure 4: Ablation study of different numbers of utility functions on “humor” vs. “helpful” and “deberta” vs. “faithful” tasks.
Variants of utility functions. We explored the impact of using linear utility functions, designed as $w^{T}x$ where $||w||_{2}=1$, and compared the performance with our original approach across all tasks. The generated Pareto fronts, as shown in Figure 5 and Appendix C.5, illustrate that our method continues to deliver strong results even with linear utility functions. However, there is still a performance gap between the results achieved with distributional Pareto-optimal utility functions and those obtained using linear functions. While linear functions still produce reasonably good trade-offs, they fail to capture the full range of optimal solutions attainable by our method, which leverages distributional utility representations. This highlights the advantage of our approach in handling more complex utility functions, demonstrating its superior flexibility in navigating multi-objective optimization.
Figure 5: Ablation study of using linear utility functions.
Constraint satisfaction and variance objective. In our case study on the “harmless” vs. “humor” task, we use 2,000 inference samples and compute both metrics for our method, Rewarded Soup, and RiC. As summarized in Table 3, our method achieves the highest overall constraint satisfaction ratio and exhibits the largest variance objective value, outperforming both Rewarded Soup and RiC. This indicates that our approach offers a more comprehensive distributional coverage, effectively handling diverse user preferences.
  Method Constraint Satisfaction Variance
RiC 0.5637 0.0662
Rewarded Soups 0.6744 0.1631
Ours 0.7184 0.5160
Table 3: Comparison of Methods for Constraint satisfaction and variance objective.
5 Conclusion
We propose Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel multi-objective LLM alignment framework. By transforming user-specified preference into symbolic tokens through a diverse ensemble of strictly increasing, non-linear utility functions, UC-MOA circumvents the numerical sensitivity issues seen in methods like RiC while avoiding the high computational costs of approaches such as MORLHF, Rewarded Soups, and MODPO. Our experiments demonstrate that UC-MOA achieves superior Pareto fronts across complex reward dimensions, evidencing robust and scalable alignment with human values.
Limitations
Representation of Human Preferences. Although UC-MOA effectively captures the distributional properties of human feedback, its performance heavily depends on the quality and diversity of the underlying preference data. Real-world human feedback is often noisy or biased, and addressing such discrepancies requires robust preprocessing and annotation methods. Additionally, certain complex preferences might not be adequately represented by the utility functions, potentially limiting the model’s applicability in diverse contexts, such as safety-critical domains like healthcare and autonomous systems.
Computational and Scaling Challenges. The proposed framework improves efficiency compared to existing methods, but it still requires substantial computational resources when scaling to tasks with numerous objectives or high-dimensional reward spaces. The reliance on large-scale GPUs and the requirement for iterative online updates may limit its adoption in settings with constrained computational budgets. Further optimization of the model architecture and training process could help mitigate these challenges, making the framework more accessible and enabling faster experimentation.
Generalizability Across Tasks. While UC-MOA has demonstrated superior alignment performance in text generation and summary tasks, its applicability to other domains, such as multimodal tasks, remains unexplored. Extending the framework to these areas might introduce additional challenges, including the need to handle dynamic environments and integrate feedback in near real-time. Adapting and validating this approach across a broader range of use cases to evaluate its generalizability and robustness is a potential direction for future research.
Ethical Considerations
While multi-objective optimization and reward modeling can enhance the alignment of large language models, our method must be applied responsibly to avoid harmful outcomes. The potential biases or harmfulness in training data or reward models may mislead the final LLM’s behavior. Careful curation of training and evaluation datasets, combined with auditing of model outputs, is essential to mitigate these risks.
Appendix A Details of Datasets
In this section, we introduce the HH-RLHF dataset and Reddit Summary dataset in more detail. We provide the links to the HH-RLHF dataset and the three reward models we used for the experiments in Table 4.
HH-RLHF Dataset
Dataset Anthropic/hh-rlhf
Harmless Reward gpt2-large-harmless-reward_model
Helpful Reward gpt2-large-helpful-reward_model
Humor Reward humor-no-humor
Reddit Summary Dataset
Dataset openai/summarize_from_feedback
Summary Reward gpt2_reward_summarization
Deberta Reward reward-model-deberta-v3-large-v2
Faithful Reward bart-faithful-summary-detector
Table 4: Datasets and Reward Models.
HH-RLHF Dataset
The HH-RLHF dataset, as described in Bai et al. (2022), was collected to support research on aligning language models using human preferences. The dataset comprises approximately 160,000 prompts each paired with two responses generated by a language model. For every prompt, one response is labeled as “chosen” (preferred by human annotators) and the other as “rejected”. Human raters evaluated paired responses based on criteria for helpfulness and harmlessness. For helpfulness, annotators selected the response that best addressed the given task, whereas for harmlessness (red-teaming) they identified the response as more likely to produce harmful content. This dual-annotation design allows for training reward models that capture subtle aspects of human judgment. Note that we follow Yang et al. (2024d) and use “chosen” data to construct our offline dataset.
Reddit Summary Dataset
The Reddit Summary dataset, introduced in Stiennon et al. (2020), is intended for research in abstractive summarization with human feedback. The dataset consists of approximately 14,900 Reddit posts paired with corresponding summaries. Each entry comprises a post and its summary, where the summary is either the original TL;DR provided by the poster or a human-annotated summary aimed at capturing the post’s main ideas.
Appendix B Implementation Details
B.1 Utility Functions
Non-Decreasing Neural Network.
We used a Non-Decreasing Neural Network to parameterize the utility function. The network architecture consisted of three fully connected layers, with our designed non-decreasing activation functions applied to the hidden layers:
$\begin{split}f(x)=\text{concat}[&max(x,-0.5),min(x,0.5),\\ &clip(x,-0.5,0.5)].\end{split}$ (8)
This design of the activation function is an extension of the ReLU activation function, which empirically makes the learned utility function more diverse, covering more types of decision boundaries. After each gradient update during training, we clip all the trainable weights in the network to be non-negative to ensure that the resulting utility function was non-decreasing. The network’s input dimension corresponded to the number of objectives in the environment, and the output dimension was a scalar representing the utility value.
Return Normalization.
We use a normalization technique by linearly mapping the return into scale $[0,1]^{K}$ without modifying the optimal policies. Specifically, we keep track of the minimum and maximum multivariate returns that the policies have currently achieved, denoted by $\mathbf{z}_{min}$ and $\mathbf{z}_{max}$. We let $\mathbf{z}_{mid}=(\mathbf{z}_{min}+\mathbf{z}_{max})/2$, and $d=\max_{i}[\mathbf{z}_{max,i}-\mathbf{z}_{min,i}]$. Before any multivariate return $z$ is fed into the utility function, we normalize it by
$z_{norm}=\frac{1}{d}(z_{norm}-z_{mid})+\frac{1}{2}$
This maps $z$ linearly to $z_{norm}$, where $z_{norm}\in[0,1]^{K}$, without affecting the scale of each dimension of returns. In the implementation, the return $\mathbf{z}_{t}$ is normalized before feeding into the utility function to compute the scalar reward.
B.2 Prompt Labeling
We provide an example of our prompt labeling technique in the “harmless” vs. “helpful” task. As illustrated in Figure 6, the response is evaluated in two dimensions “harmless” and “helpful”. The two-dimensional scores are further feed into ten utility functions and obtain a set of utility scores. We then convert these scores into percentile rankings. Finally, we label the prompt with the utility function whose percentile ranking is highest. In the example, the highest percentile ranking belongs to utility index 2 (67.67%), so we append the prompt with <max_utility_idx> c. This procedure ensures that each prompt-response pair is labeled with the most prominent utility dimension, i.e., the one where its performance stands out most relative to other examples, thereby facilitating more focused downstream alignment.
Figure 6: An example of our prompt labeling technique in “harmless” vs. “helpful” dataset.
B.3 Hyperparameter Choices
Table 5 summarizes the hyper-parameter setting for our experiments.
Training Phase
Offline Learning Rate $1.414\times 10^{-4}$
Online Learning Rate $10^{-5}$
Batch Size assistant: $16$
summary: $2$
Online Generation kwargs max_new_tokens (assistant): $128$
max_new_tokens (summary): $48$
num_of_samples: $1000$
min_length: $-1$
top_k: 00
top_p: $0.9$
temperature: $0.9$
do_sample: True
begin_suppress_tokens: [tokenizer.eos_token_id]
Evaluation Phase
Generation kwargs max_new_tokens (assistant): $128$
max_new_tokens (summary): $48$
min_length: $-1$
top_k: $0.0$
top_p: $0.9$
do_sample: True
Table 5: Hyper-parameter settings in our experiments.
B.4 Distributional Metrics
We use two novel metrics as in Cai et al. (2023) for assessing distributional preference: Constraint Satisfaction and Variance Objective.
For Constraint Satisfaction, we define $u(p,\pi)$ to be the probability of return random variance $z\sim\mathbf{Z}(\pi)$ that satisfies the constraint $p$, representing how a policy satisfies the given constraint. We randomly generate each constraint defined by $n$ linear constraints represented by $\{\mathbf{z}|\forall i,\mathbf{w}_{i}^{T}\mathbf{z}\geq c_{i}\}$, where $w_{i}$ is uniformly sampled from the surface $\mathbf{w}_{i}\succeq 0,\mathbf{1}^{T}\mathbf{w}_{i}=1$, and $c_{i}$ is sampled from the distribution of $\mathbf{U}(\min_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z},\max_{\mathbf{z}\in\mathbf{Z}}\mathbf{w}_{i}^{T}\mathbf{z})$, where $\mathbf{Z}$ is the set of all return samples of the final policies learned by DPMORL.
For Variance Objective, we define $u(p,\pi)$ to be a weighted sum of expected return $\mathbb{E}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}\mathbf{z}$ and standard deviation of return $\sqrt{\text{Var}_{\mathbf{z}\sim\mathbf{\mu}(\pi)}[\mathbf{z}]}$. We randomly generate $\mathbf{w}_{i}\in\mathbb{R}^{2K}$, where $\mathbf{w}_{i}\succeq 0$ and $\mathbf{1}^{T}\mathbf{w}_{i}=1$. We let $\mathbf{w}_{i}=[\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}]$, where $\mathbf{w}_{i}^{1},\mathbf{w}_{i}^{2}\in\mathbb{R}^{K}$, and use $\mathbf{w}_{i}^{1}$ and -$\mathbf{w}_{i}^{2}$ as the weight for expected return and standard derivation of return respectively.
For both of the metrics, we define $u(p,\pi)$ to be the satisfaction score of policy $\pi$ under preference $p$. We generate $M=100$ number of preference $p_{1},p_{2},\cdots,p_{M}$. For both of the Constraint Satisfaction and Variance Objective metric, the mean score is defined as
$\frac{1}{M}\sum_{i=1}^{M}\max_{j\in\{1,2,\cdots,N\}}u(p_{i},\pi_{j})$
The mean score in the above equation measures the best performance of satisfying the preference of the user, averaged in $M=100$ preferences. A higher mean score represents a better ability to meet different preferences.
Appendix C Additional Experiment Results
C.1 Experiment I
In addition to the main results presented in the paper, Figure 7 shows the Pareto fronts for the remaining four tasks in our experiments. We compare five methods: Llama 2 base, SFT, Rewarded Soups, RiC, and our proposed approach (Ours). Each subplot corresponds to a different two-objective trade-off, where the $x$- and $y$-axes represent the average test rewards for two distinct objectives. A method demonstrating an outer Pareto curve indicates a more favorable performance, as it captures a broader spectrum of trade-offs between competing objectives. Across all four subplots in Figure 7, Our method (purple line) achieves the most outward Pareto front, reflecting superior coverage of the objective space compared to the other baselines. RiC (orange line) also performs strongly but is dominated by Ours in several regions of the objective space. Meanwhile, Rewarded Soups (red line), SFT (green crosses), and Llama 2 base (blue triangles) display varying degrees of suboptimality in balancing the two objectives simultaneously. Overall, these results reinforce our main findings, showing that our method provides a robust and effective means of navigating multi-objective trade-offs across a range of different tasks.
Figure 7: Pareto fronts of the remaining four tasks.
C.2 Experiment III
We investigate the impact of the online training stage by comparing three variants of our method: (1) pure offline training (w/o online), (2) one online iteration (w/ online iter 1), and (3) two online iterations (w/ online iter 2). As shown in Figure 8, each subplot corresponds to the Pareto fronts for a different task. Across all four tasks, w/o online (blue line) consistently produces the weakest Pareto front, indicating suboptimal performance when balancing multiple objectives. Introducing a single online iteration (w/ online iter 1, green line) noticeably expands the Pareto front, suggesting that even a brief online training phase can significantly improve multi-reward optimization. Finally, adding a second online iteration (w/ online iter 2, red line) further refines these trade-offs, achieving the most outward (and thus superior) Pareto front among the three variants. These results confirm the necessity of the online stage, demonstrating that multiple online iterations can better align the model with the target objectives.
Figure 8: Ablation study of the online stage on the remaining four tasks.
C.3 Experiment IV
We examine how closely our model’s generated responses align with a user-specified target objective, in this case, the 9th utility function (index 8). Figure 9 shows the distribution of normalized utility values across all ten reward models for the “harmless” vs. “humor” task, and Table 6 reports the corresponding mean normalized values. Notably, the 9th utility function (index 8) attains the highest average score of 0.68, demonstrating that the LLM effectively prioritizes the user-indicated objective. In other words, when prompts include the requirement <max_utility_index> i specifying the 9th utility, the model consistently generates responses that yield higher scores along that dimension than along the others. These findings confirm that our fine-tuned model can reliably follow the user-designated target objective during inference.
Figure 9: Utility value distribution for the “harmless” vs. “humor” task. The required utility index is 8 (i.e., the 9th utility function).
Utility index 0 1 2 3 4 5 6 7 8 9
Mean 0.57 0.57 0.56 0.56 0.57 0.57 0.49 0.55 0.68 0.58
Table 6: Mean normalized value for each utility function.
C.4 Experiment V
We investigate how varying the number of utility functions (5, 10, 15) affects our method’s ability to balance multiple objectives. Figure 10 shows the resulting Pareto fronts across four different tasks, where an outward-shifted curve indicates better overall performance. We observe that using only 5 utility functions (blue line) yields a slightly weaker Pareto front compared to the other two settings. In contrast, the results for 10 (green line) and 15 (red line) utility functions are nearly identical, suggesting that increasing beyond 10 utility functions does not lead to a substantial performance gain. Consequently, our method demonstrates robustness to the exact number of utility functions, so long as it is sufficiently large to capture the relevant trade-offs among the objectives.
Figure 10: Ablation study of different numbers of utility functions on remaining four tasks.
C.5 Experiment VI
Figure 11: Abalation study of using linear utility functions.
We compare the performance of our method when using linear utility functions (Ours (linear)) versus our original, distribution-based approach (Ours (param)) in Figure 11. In each subplot, the curves depict how effectively each method covers the trade-offs between two objectives. Although the linear approach maintains a reasonable Pareto front and thus remains effective for balancing multiple rewards, it consistently trails the distribution-based version in all tasks. This shortfall indicates that purely linear utility functions, while simpler to construct, capture a narrower range of potential trade-offs and consequently yield a less comprehensive frontier. In contrast, our distribution-based design is more flexible, enabling better alignment with diverse user preferences and resulting in a noticeably more outward Pareto front.